OpenStudy (anonymous):

Does the following equation have a solution? Explain why or why not, and support your answer with a graph: abs(-x^2-3x+5)=-4

OpenStudy (anonymous):

it is unlikely that the absolute value of anything is negative, since the absolute value of anything is always greater than or equal to zero

OpenStudy (anonymous):

When the absolute value is presented like this in the equation it can be a negative - if you wanted to get rid of the absolute value portion for solving reasons, you would just take the positive of everything inside the absolute value brackets and continue.

OpenStudy (anonymous):

is this the question \[|-x^2-3x+5|=-4\]??

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

then there is no solution because the expression on the left is never negative, but \(-4\) is negative

OpenStudy (anonymous):

But wouldn't we be able to solve it for x? Cause I can get to \[x^2+3x=1\] but I just don't know what to do from there.

OpenStudy (blank ):

I want to reassure you that it can be solved for x. http://www.wolframalpha.com/input/?i=%28-x%5E2-3x%2B5%29%3D-4

OpenStudy (anonymous):

Sweet, thanks!

OpenStudy (anonymous):

Wait. It's true that \[x^2 + 3x = 1\] has a solution

OpenStudy (anonymous):

However, your first equation does not

OpenStudy (anonymous):

Absolute value implies that both (-x^2-3x+5)=-4 and (-x^2-3x+5)=4 have a solution

OpenStudy (anonymous):

(-x^2-3x+5)=4 has a solution, but it's complex

OpenStudy (anonymous):

That complex solution would have square roots and fractions, right?

OpenStudy (anonymous):

It's square root of a negative number. The solution is (-3 +- (rad)-4)2

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%28-x%5E2-3x%2B5%29%3D-4 because this says it can be solved, I just can't get there